DAVENPORT PAIRS OVER FINITE FIELDS

WAYNE AITKEN MICHAEL D. FRIED LINDA M. HOLT

Abstract. A pair of polynomials f, g ∈ 𝔽_{q}[T] is called a Davenport pair if their value sets are equal, V_{f}(𝔽_{q^{t}}) = V_{g}(𝔽_{q^{t}}), for infinitely many extensions of 𝔽_{q}. This paper studies such pairs.